# Lepton seesaw model in a modular $A_{4}$ symmetry

### v1

@@ III Numerical analysis and phenomenology
Figure 1: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where the blue color represents the range of $0-1$, the green $1-2$, the yellow $2-3$, and the red $3-5$ $\sigma$ standard deviation estimated by $\sqrt{\Delta\chi^{2}}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. The central figure represents the allowed range of Majorana phases. The right figure demonstrates the allowed Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound on the DESI and CMB data.
III.1 NH
@@
In this subsection, we discuss the allowed regions in case of NH. In Figs. 1, the blue color represents the range of $0-1$, the green $1-2$, the yellow $2-3$, and the red $3-5$ $\sigma$ standard deviation estimated by $\sqrt{\Delta\chi^{2}}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. It suggests that we have enough solutions at nearby $\tau=\omega$. The central figure represents the allowed range of Majorana phases $\alpha_{21}$ and $\alpha_{31}$. We find $\alpha_{21}$ is almost zero degree but $\alpha_{31}$ is allowed by the ranges [$177.5^{\circ}-178.5^{\circ}$] and [$181^{\circ}-182.5^{\circ}$].
@@
The right figure of Fig. 1 demonstrates the allowed region of Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound from the combination of the DESI and CMB data. $\delta_{\rm CP}$ is allowed by the ranges of [$80^{\circ}-100^{\circ}$] and [$240^{\circ}-300^{\circ}$]. $\sum D_{\nu}$ is allowed by the range of [$0.22\ {\rm eV}-0.35\ {\rm eV}$].
@@
Figure 2: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV.
@@
In Fig. 2, we show some predictions in the case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. The red and gray dotted lines respectively correspond to upper and lower bounds of $\langle m_{ee}\rangle$ experiment at 90 % confidence level. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.07-0.12]$ [eV] that is completely within the range of the experiment. The lightest neutrino mass goes over the range of $D_{\nu_{1}}=[0.07-0.11]$ [eV] whose range is similar to the one of neutrinoless double beta decay. Even though our $\sum D_{\nu}$ does not satisfy the cosmological bound $0.12$ [eV], $m_{\nu_{e}}^{2}$ is totally safe to the experimental upper bound.
III.2 IH
@@
The right figure of Fig. 3 demonstrates that $\delta_{\rm CP}$ is allowed by the ranges of [$260^{\circ}-280^{\circ}$] while $\sum D_{\nu}$ is allowed by the range of [$0.44\ {\rm eV}-0.50\ {\rm eV}$] that is a typical tendency.
@@
Figure 4: Numerical $\Delta\chi^{2}$ analyses in case of IH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV.
@@
In Fig. 4, we show the same kinds of figures as in Figs. 2, and color legends and the captions are the same as the ones of Fig. 2. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.14-0.16]$ [eV] that is nearby the range of upper bound of this experiment. Therefore, this case has high testability that could be confirmed/excluded in near future. The lightest neutrino mass goes over the range of $D_{\nu_{3}}=[0.14-0.16]$ [eV] whose range is similar to the one of neutrinoless double beta decay. The tendency is likely to the case of NH. $m_{\nu_{e}}^{2}$ has the same tendency as the case of NH. Therefore, it is totally safe to the experimental upper bound.
IV Summary and discussion
@@
We have proposed a lepton seesaw model in a modular $A_{4}$ symmetry. The charged-lepton mass matrix and the neutrino mass matrix have the same origin associated with vector-like fermions. As a result, the charged-lepton mass matrix is induced via seesaw-like mechanism, while the neutrino mass matrix is generated through the inverse seesaw mechanism where the tiny Majorana mass matrix is supposed to be obtained via the VEV of $\Delta$.
@@
In addition we have searched for allowed regions at nearby fixed points which are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021), and we have found the solutions at nearby $\tau=\omega$. In our chi square numerical analysis, we have found large value of sum of the neutrino masses which exceed the cosmological upper bound $0.12$ eV for both the cases of NH and IH. However, our solutions are totally safe for the direct search for neutrino mass which is done by the Karlsruhe Tritium Neutrino experiment for both the cases. In addition, our region of neutrinoless double beta decay is localized at nearby the experimental bound, and therefore it would be tested near future. We have obtained narrow ranges of Majorana and CP phases which would also be verifiable via some experiments.
@@
Finally, we would like to mention our next task. Even though we have successfully built our lepton seesaw model which has the same mass origin from vector-like fermions, our model has an assumption among mass hierarchies such as $m^{\prime}\ll M_{L^{\prime}}$. This assumption is in a sense very important to construct the seesaw model. Thus, we will move to a model that would have appropriate and natural mass hierarchies among these mass matrices.
@@ Acknowledgments
The work was supported by the Fundamental Research Funds for the Central Universities (T. N.).

### v2

@@ IV Numerical analysis and phenomenology
Figure 1: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where the blue color represents the range of 010-1, the green 121-2, the yellow 232-3, and the red 353-5 $\sigma$ standard deviation estimated by $\sqrt{\Delta\chi^{2}}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. The central figure represents the allowed range of Majorana phases. The right figure demonstrates the allowed Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound on the DESI and CMB data.
IV.1 NH
@@
In this subsection, we discuss the allowed regions in case of NH. In Figs. 1, the blue color represents the range of 010-1, the green 121-2, the yellow 232-3, and the red 353-5 $\sigma$ standard deviation estimated by $\Delta\chi^{2}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. It suggests that we have enough solutions at nearby $\tau=\omega$. The central figure represents the allowed range of Majorana phases $\alpha_{21}$ and $\alpha_{31}$. We find $\alpha_{21}$ is almost zero degree but $\alpha_{31}$ is allowed by the ranges $177.5^{\circ}-178.5^{\circ}$] and $181^{\circ}-182.5^{\circ}$].
@@
The right figure of Fig. 1 demonstrates the allowed region of Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound from the combination of the DESI and CMB data. $\delta_{\rm CP}$ is allowed by the ranges of $80^{\circ}-100^{\circ}$] and $240^{\circ}-300^{\circ}$]. $\sum D_{\nu}$ is allowed by the range of $0.22\ {\rm eV}-0.35\ {\rm eV}$].
@@
Figure 2: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV. Cyan region is allowed by experimental result of Nufit 6.0.
@@
In Fig. 2, we show some predictions in the case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. The red and gray dotted lines respectively correspond to upper and lower bounds of $\langle m_{ee}\rangle$ experiment at 90 % confidence level. Cyan region is allowed by experimental result of Nufit 6.0. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.07\ {\rm eV}-0.12\ {\rm eV}]$ that is completely within the range of the experiment. The lightest neutrino mass spans over the range of $D_{\nu_{1}}=[0.07\ {\rm eV}-0.11\ {\rm eV}]$ whose range is similar to the one of neutrinoless double beta decay. Even though our $\sum D_{\nu}$ does not satisfy the cosmological bound $0.12$ [eV], $m_{\nu_{e}}^{2}$ is totally safe to the experimental upper bound.
parameter BF parameter BF parameter BF
$\tau$ 0.503+0.916i-0.503+0.916i $\tilde{y}_{\Delta}$ 0.000582005+0.00028466i-0.000582005+0.00028466i $\tilde{f}_{2}$ $51.7919$
$\tilde{f}_{3}$ 8.42753 $\tilde{M}_{2}$ $42.6753$ $\tilde{M}_{3}$ $0.00620328$
$\tilde{a}_{e}/{\rm GeV}$ $0.000361577$ $\tilde{b}_{e}/{\rm GeV}$ $0.0761912$ $\tilde{c}_{e}/{\rm GeV}$ $1.29171$
$s_{12}$ 0.547201 $s_{23}$ 0.682414 $s_{13}$ 0.150202
$\Delta m^{2}_{\rm sol}$ $7.45632\times 10^{-5}~{\rm eV}^{2}$ $\Delta m^{2}_{\rm atm}$ $2.52784\times 10^{-3}~{\rm eV}^{2}$ $\delta_{\rm CP}$ $73.7719^{\circ}$
$\alpha_{21}$ $1.19252^{\circ}$ $\alpha_{31}$ $181.673^{\circ}$ $\kappa$ $48.7~{\rm meV}$
$m^{2}_{\nu e}$ $0.00776348~{\rm eV}^{2}$ $\sum D_{\nu}$ $276.805~{\rm meV}$ $\langle m_{ee}\rangle$ $87.7089~{\rm meV}$
Table 2: Best-fit (BF) parameter values in the NH case corresponding to $\Delta\chi_{\rm min}=1.56$, where $\tilde{a}_{e}\equiv\frac{vv^{\prime}}{2}\frac{a_{e}f_{1}}{M_{1}}$, $\tilde{b}_{e}\equiv\frac{vv^{\prime}}{2}\frac{b_{e}f_{2}}{M_{2}}$, and $\tilde{c}_{e}\equiv\frac{vv^{\prime}}{2}\frac{c_{e}f_{3}}{M_{3}}$.
IV.2 IH
@@
The right figure of Fig. 3 demonstrates that $\delta_{\rm CP}$ is allowed by the ranges of $260^{\circ}-280^{\circ}$] while $\sum D_{\nu}$ is allowed by the range of $0.44\ {\rm eV}-0.50\ {\rm eV}$] that is a typical tendency.
@@
Figure 4: Numerical $\Delta\chi^{2}$ analyses in case of IH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV. Cyan region is allowed by experimental result of Nufit 6.0.
@@
In Fig. 4, we show the same kinds of figures as in Fig. 2, and color legends and the captions are the same as the ones of Fig. 2. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.14\ {\rm eV}-0.16\ {\rm eV}]$ that is nearby the range of upper bound of this experiment. Therefore, this case has high testability that could be confirmed/excluded in near future. The lightest neutrino mass spans over the range of $D_{\nu_{3}}=[0.14\ {\rm eV}-0.16\ {\rm eV}]$ whose range is similar to the one of neutrinoless double beta decay. The tendency is likely to the case of NH. $m_{\nu_{e}}^{2}$ has the same tendency as the case of NH. Therefore, it is totally safe to the experimental upper bound.
parameter BF parameter BF parameter BF
$\tau$ 0.516664+0.874007i-0.516664+0.874007i $\tilde{y}_{\Delta}$ 0.0492088+0.0215522i-0.0492088+0.0215522i $\tilde{f}_{2}$ $0.213211$
$\tilde{f}_{3}$ 8.24262 $\tilde{M}_{2}$ $15.2601$ $\tilde{M}_{3}$ $0.00635192$
$\tilde{a}_{e}/{\rm GeV}$ $0.00034419$ $\tilde{b}_{e}/{\rm GeV}$ $0.0726524$ $\tilde{c}_{e}/{\rm GeV}$ $1.23426$
$s_{12}$ 0.558474 $s_{23}$ 0.71674 $s_{13}$ 0.1497
$\Delta m^{2}_{\rm sol}$ $7.47522\times 10^{-5}~{\rm eV}^{2}$ $\Delta m^{2}_{\rm atm}$ $2.53455\times 10^{-3}~{\rm eV}^{2}$ $\delta_{\rm CP}$ $277.664^{\circ}$
$\alpha_{21}$ $0.442^{\circ}$ $\alpha_{31}$ $180^{\circ}$ $\kappa$ $76.3653~{\rm meV}$
$m^{2}_{\nu e}$ $0.023038~{\rm eV}^{2}$ $\sum D_{\nu}$ $447.588~{\rm meV}$ $\langle m_{ee}\rangle$ $151.663~{\rm meV}$
Table 3: Best-fit (BF) parameter values in the IH case corresponding to $\Delta\chi_{\rm min}=2.14$, where $\tilde{a}_{e}\equiv\frac{vv^{\prime}}{2}\frac{a_{e}f_{1}}{M_{1}}$, $\tilde{b}_{e}\equiv\frac{vv^{\prime}}{2}\frac{b_{e}f_{2}}{M_{2}}$, and $\tilde{c}_{e}\equiv\frac{vv^{\prime}}{2}\frac{c_{e}f_{3}}{M_{3}}$.
V Summary and discussion
@@
We have proposed a lepton seesaw model in a modular $A_{4}$ symmetry. The charged-lepton mass matrix and the neutrino mass matrix have the same origin associated with vector-like fermions. As a result, the charged-lepton mass matrix is induced via seesaw-like mechanism, while the neutrino mass matrix is generated through the inverse seesaw mechanism where the tiny Majorana mass matrix is supposed to be obtained via VEV of $\Delta$.
@@
In addition we have searched for allowed regions at nearby fixed points which are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021), and we have found the solutions at nearby $\tau=\omega$. In our chi square numerical analysis, we have found large value of sum of the neutrino masses which exceed the cosmological upper bound $0.12$ eV for both the cases of NH and IH. However, our solutions are totally safe for the direct search for neutrino mass which is done by the Karlsruhe Tritium Neutrino experiment for both the cases. Our region of neutrinoless double beta decay is localized at nearby the experimental bound, therefore it would be tested near future. We have obtained narrow ranges of Majorana and CP phases which would also be verifiable via some experiments.
@@
Finally, we would like to mention our next task. Even though we have successfully built our lepton seesaw model which has the same mass origin from vector-like fermions, our model has an assumption among mass hierarchies such as $m^{\prime}\ll M_{L^{\prime}}$. This assumption is very important to construct the seesaw model buildings. Thus, we will move to a model that would have appropriate and natural mass hierarchies among these mass matrices.
@@ Acknowledgments
The project is supported by the Fundamental Research Funds for the Central Universities (T. N.) and Zhongyuan Talent (Talent Recruitment Series) Foreign Experts Project (H. O.).